Please invent a random variable (be sure it follows the rule that the probabilities must add to 1). Or find a random variable in the text. Then find E(X) showing all the steps.
Answer #1

Answer :. Consider a random variable x that can take value based on the number that comes on dice.

So x = 1,2,3,4,5,6

P(1) = p(2) = p(3) = p(4) = p(5) = p(6) = 1/6

So expected value of x :

E[x] = summation ( pi * xi) for all values I of x

= (1*1/6 + 2*1/6 +3*1/6 +4*1/6 +5*1/6 + 6*1/6) =21/6 = 7/2